A bag contains 3 pink gumballs, 4 red gumballs, and 2 green gumballs. Find the probability that you do not draw all green gumballs if there are 3 drawings that occur and the gumballs are replaced after each draw. Round your answer to two decimal places.
Answer #1

Total = 3 + 4 + 2 = 9

Green = 2

P(not all 3 are green) = 1 - P(all three green) = 1 - (2/9)*(2/9)*(2/9) = 721/729 = 0.99

Hence, 0.99 is the probability that not all 3 drawings will be green.

I hope this will help. Thank you.